The life expectancy at birth of both sexes in a particular year can be described by the function f(x)=11.36+14.96 In x, where x represents the number of years after 1900. Use this function to determine the life expectancy in 2036.
Answer by stanbon(75887)

The life expectancy at birth of both sexes in a particular year can be described by the function f(x)=11.36+14.96 In x, where x represents the number of years after 1900. Use this function to determine the life expectancy in 2036.
-------
f(36) = 11.36 + 14.96*ln(36) = 65 when rounded to a whole number.
Cheers,
Stan H.
---------